OpenStudy (darkbluechocobo):

Calculate the elastic potential energy of a spring with a spring constant of 50N/m and a distance of (x) of 20cm

OpenStudy (shamim):

Given k=50N/m x=20cm=0.2m P.E=?

OpenStudy (shamim):

Required formula is P.E=(1/2)*k*x^2=?

OpenStudy (darkbluechocobo):

25*.04

OpenStudy (darkbluechocobo):

P.E.=1? @Shamim

OpenStudy (shamim):

Ys u r right